OpenStudy (anonymous):

Need help on these two questions on my math review. A car moving with a speed of 108 km/h is brought to rest in a distance of 60 m. How much time did the car take to stop? 2 seconds 4 seconds 6 seconds 8 seconds An object is thrown with an initial velocity of 3 m/s. What is its acceleration at its maximum height? zero 3 m/s2 -9.8 m/s2 none of the above

OpenStudy (owen3):

can we assume constant acceleration?

OpenStudy (owen3):

or deceleration in this instance

OpenStudy (anonymous):

yes?

OpenStudy (owen3):

there are equations we can use x = 1/2 * a * t^2 + v*t v^2 = u^2 + 2*a*s

OpenStudy (owen3):

1 km = 1000 m 1 m = 1/1000 km v^2 = u^2 + 2*a*s 60 m = 60* 1/1000 km 0^2 = 108^2 + 2*a*60*1/1000 a = -108^2 / ( 2 * 60 /1000 )

OpenStudy (owen3):

a = -97200 now plug this into v = u + a*t 0 = 108 + (-97200) * t -108 / -97200 = t note that time is in hours to get seconds multiply by 3600

OpenStudy (owen3):

The next question, the acceleration due to gravity is always -9.8 , near the surface of earth.

OpenStudy (anonymous):

-9.8 m/s^2